Startup Night 2: 2017 Sep 03 UT

Observers: OKuhn, BRothberg (Tucson) SWalsh (LBTO)
Telescope Operator: SAllanson (LBTO)
AO Support: GTaylor
Telescope Support: KNewton (LBTO)
Software Support: CBiddick, KSummers (Tucson)
Instrument Support: ESolheid (LBTO)
Engineering Support: TMcMahon

Summary

Overcast and no stars visible in the all-sky, but it started to break up a few hours before sunrise. Shane carried out LUCI flexure compensation calibrations ( https://docs.google.com/document/d/17-jxVB7P48NaTkUP5qJhnxOk9RJVUo8jot3yJAK9Up4/edit?ts=59ab5fbb#) until we opened at ~03:30.

During LUCI cals, ran LBCs in the closed dome (mostly dark) for further checkout:

Tracker images look OK (taken in dark dome but all 4 look uniform and show no pickup noise or banding)

180-s exposures in dark dome: odd scattered light effects seen for LBCR V-Bessell, R-Bessell and r-SLOAN. Prompted a check of the filters which all look pretty dirty - moths already have soiled these filters. One shows a drip.

LBCR biases and these low-light-level 'flats' show intermittent horizontal banding (IT #6748). I did not see this banding in the on-sky images taken last night, but so far only in biases and low-light 'dome flats'. Amplitude of effect is small, ~1% or 5 ADU above a ~500 ADU background.

Ran an OB on NGC628 @ Us and r-SLOAN to test guiding and dithering. Images are guided and dithered. Tech chip focus corrections are computed (not yet verified that they are sent) as I see them in the lbc.log file. Not very good IQ initially on both blue and red, but seemed a bit better after refocus.

[Odd behavior of LBC's resume from pause. It's not widely used, so perhaps I don't remember how it works. Click play after pause and resumes at next dither, as expected, but TCS receives this as a preset (not just an offset).]

Tried the NS target again, but at 11:50 so there was not a broad time span left to cover. May wish to retry. We also noticed that, unlike last night, the telescope did not dither (ran *V2.xml). Even though offsets were being sent to PCS, they were not always propagated. Steve toggled off/on the PCS flag for varying rates but still this did not work. End-of-night mystery...

Sky flats at B and I (at 0.3-sec and one PA).

Preparations

Kellee got both LBC trackers running. Blue tech PC slow to boot, and is now configured to remain on when you turn off LBC systems. If you must power cycle blue tech, you can do this through power.py.

The 4D laser was found to be on and turned off. Most likely this was the cause of the speckling reported in IT #6744. No sign of speckling tonight, not on long, 300-sec, closed dome exposures or on ~60-sec sky exposures.

LBCs came up on first try. There is some horizontal banding in LBCR biases and low-light level images (180-sec exposures in a dark dome while LUCI calibrations are on-going and LBCs are out of the beam... so one might expect weird scattering effects on images).

Log

04:15 we noticed some "drips" on LBCR V image. Looking back, these are on earlier V, and also R and r images. Also at 04:15 the LBCR hardware failure on filter wheel #2. Steve brought telescope to horizon and he and Shane are going up to have a look. Moths! Already there is dust and even a drip on the filters: 2' (22 R-BESSEL) is the dustiest. 3' (23 V-BESSEL has dust. 1 (15 i-sloan) has a drip.

05:15 Reconfiguring for LBCs before we go back to more LUCI flex comp calibrations.

10:15 Last flex comp script will finish soon. Clearing up so the plan is to use LBC. Meanwhile, noticed scallop along left edges of LBCB chips (see lbcb 083924) and reviewed IT on this #3598 where it was concluded either to be florescence or flat field pattern.

10:39 Opening up.

10:55 Slew to copoint for NGC628 field on which we will test guiding, tech chip focus and dithering. 5-pt dither at Us and r. Looks clear on allsky.

10:56 dohybrid. (Similar to last night as I recall - init blue pupils too big, red, too small. Tamb=11.6 C). Converged in 2 fpia iterations after wrs.

11:04 Copoint. b 110440 and r 110435

MOUNT: delta_IE = -11.12 arcsec, delta_CA = -2.14 arcsec
M1 SX: dX = 0.66 mm, dY = -0.05 mm, dRX = -0.98 arcsec, dRY = -14.21 arcsec
M1 DX: dX = -0.68 mm, dY = 0.08 mm, dRX = 1.79 arcsec, dRY = 14.55 arcsec

lbcrangebal got confused, it appeared to work, as if it had displayed an image, but I could not find it. Then I tried to start it again, getting the variable not found error. Restarted both iraf & ds9 finally.

11:12: Run NGC628 OB Us 160-s and r-SLOAN 2x60-sec.Guiding working. ActOp corrections appear to be computed on Red side. On blue, pupils must just be too faint. IQ on blue side is not good.

11:20 Stop OB to run dofpia once more to see whether IQ can be improved.The last blue and red images are both bad. See b 111422 r 111747. Glass and ambient are within 0.6 and 0.8 deg. Converged in 4 iterations.

11:27 Run NGC628 again. Blue guide star is less elongated, so it seems refocus improved things. Seeing ~0.9-1" Blue images look better. Red images show some coma near top of chip2. After 2 blue exposures, I paused the OB. I decided to take one more blue exposure (1 more dither position) but I believe that when I clicked play, the OB just started from the beginning (Steve received a preset, not just an offset, however I see what I expect - that the OB continued on to the 3rd dither position).

lbcb.20170903.112836.fits NGC628 0 0
lbcb.20170903.113207.fits NGC628 -20 -40
lbcb.20170903.113649.fits NGC628 20 40

LBC Non-sidereal tests

11:42 Al's target: dofpia. 4 iter

11:48 Slew to NS target (hijacked). Unticked guiding. VV 60-sec. OSIRIS_Rex_LBC_V2.xml

12:00 Although the OB should dither, the images are not dithered. PCS showed the first two offsets correctly, but the third and the others were 0, 0. And although PCS did display non-zero offsets at least once, the images are not dithered. The third image 120533 shows a jump of ~60 arcsec midway through the image. Dithering did work for the sidereal target (NGC628). And dithering worked on this field yesterday, so I am not sure why it is not now.

lbcr.20170903.120159.fits
lbcr.20170903.120346.fits
lbcr.20170903.120533.fits
lbcr.20170903.120720.fits
lbcr.20170903.120909.fits

Try again with OB reloaded, played from beginning.

Steve disabled and reenabled the PCS flag which deals with varying rates and we'll try again. It's getting light, so I had to decrease the exptime by a factor 10 for this test. Seems offsets are not what are expected and the field itself, regardless of what PCS said, shows no offsets from image to image. A mystery...

12:31 BI flats with 0.3 sec exposure times.

 
Topic revision: r3 - 03 Sep 2017, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback